• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 19:19:41 on March 25, 2009

    Entry number 266555

    Followups:

    keyword=L-HRS Delay Removed for S1 and S2m TDC Signals

    I bypassed the PEC delay boxes for S1 and S2m on the left HRS. This was
    necessary, since we are now using a common stop TDC (1877) instead of the
    1875. I haven't been able to check my changes, since there is no L1A
    signal coming from the BigBite weldment. Looking at the BigBite
    trigger supervisor, it appears most of the triggers are disconnected.
    This would explain the lack of an L1A signal.

    I needed to send the S2m left and right PMT signals into two LeCroy 4518
    modules so that I send a copy of the signals to the scalers. I used two
    modules that had been used for the A1 Aerogel during Transversity. The
    S2m left signals are in slot 11, and the S2m right signals are in slot 14.